# A Multi-Perspective Analysis of Memorization in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs), trained on massive corpora with billions of parameters, show unprecedented performance in various fields. Though surprised by their excellent performances, researchers also noticed some special behaviors of those LLMs. One of those behaviors is memorization, in which LLMs can generate the same content used to train them. Though previous research has discussed memorization, the memorization of LLMs still lacks explanation, especially the cause of memorization and the dynamics of generating them. In this research, we comprehensively discussed memorization from various perspectives and extended the discussion scope to not only just the memorized content but also less and unmemorized content. Through various studies, we found that: (1) Through experiments, we revealed the relation of memorization between model size, continuation size, and context size. Further, we showed how unmemorized sentences transition to memorized sentences. (2) Through embedding analysis, we showed the distribution and decoding dynamics across model size in embedding space for sentences with different memorization scores. The n-gram statistics analysis presents d (3) An analysis over n-gram and entropy decoding dynamics discovered a boundary effect when the model starts to generate memorized sentences or unmemorized sentences. (4)We trained a Transformer model to predict the memorization of different models, showing that it is possible to predict memorizations by context.
@@ 1 Introduction
Large language models (LLMs), spanning from BERT Devlin et al. (2019) to the recent strongest GPT-4 OpenAI et al. (2024), have not just revolutionized the Natural Language Processing (NLP) but nearly the whole artificial intelligence research field. Though their performances on various tasks are surprising, we still know little about the underlying mechanism of Large Language Models whose structures are based on neural models known as black-box Alain and Bengio (2018). Due to the enormous model size and pre-train data size, unprecedented phenomena in previous machine learning are also identified in the LLM, while in this research we discuss the memorization.
Memorization Hartmann et al. (2023) in the context of LLMs means the LLM sometimes will verbatim output the same sentence recorded in their pre-training corpus. Memorization is a coin with two sides, on the one hand, utilizing such memorization, we can use LLMs as a knowledge base Petroni et al. (2019) that can be served for multiple usage, on the other hand, personal information contained in the pre-train corpora may also be elicited maliciously Yao et al. (2024). Though such a phenomenon has been utilized or studied since the merge of LLMs, detailed analysis of memorization is still under-explored.
@@ 2.1 Scaling Laws of LLM
In this research, the conducted experiments span across various model sizes, thus relating to the research of Scaling Laws Kaplan et al. (2020); Abnar et al. (2021); Villalobos (2023), which suggests the LLM’s performance scales with the size of corpora, the parameter of model and the computation required.
Those researches inspired LLMs to continuously scale those factors to gain higher performance like T5 Raffel et al. (2023), GPT-3 Brown et al. (2020), PaLM 2 Anil et al. (2023), etc. On the other hand, researchers also analyzed how LLM Scaling affects particular tasks rather than general performance. For example, Sun and Miceli-Barone (2024) discussed how LLM scaling affects both translation tasks and prompt injection attacks. Within those discussions, the LLM’s emergent abilities Wei et al. (2022) are also discovered, which means the LLM suddenly reaches high performance on previous low-performance tasks when the LLM reaches a specific model size such as digital addition Nye et al. (2021), reasoning Lampinen et al. (2022) etc. Recent studies have questioned whether emergent abilities are just mirages caused by misuse of metrics Schaeffer et al. (2023) or whether emergent abilities are just context-learning Lu et al. (2023).
@@ 2.2 Memorization
Memorization, which means the neural model tends to memorize the case during training, is a long-existing phenomenon in machine learning. While the emergence of LLMs recalled this topic.
Before LLM, a phenomenon close to memorization in real-world practices is known as over-fitting Tetko et al. (1995) which means a near-zero loss in the train set while bad performance in test set. Zhang et al. (2017) described that the neural model could reach near-zero loss on the training set, which suggests the neural model perfectly memorized the relation between input and data label. However, memorization is different from over-fitting as LLM has good performance in the test set or the real world. Feldman (2020) has conducted a theoretical analysis explaining why memorization happened and is needed. They showed in a long-tail distribution where many categories only have very few data. In this situation, as the number of data does not support extracting general features to describe the long-tail data, the best choice for the neural model is just to memorize them and compare them with data in the test set.
In the era of LLM, which differs from previous classification models, the LLM can directly output their pre-training data also known as memorization, previous researchers extracted knowledge bases from LLM Petroni et al. (2019); AlKhamissi et al. (2022) showing massive knowledge memorized by LLM. However, the massive knowledge learned during pre-training also leads to the concern of data contamination Sainz et al. (2023) and privacy risk Yao et al. (2024). Regarding specific research of memorization behavior itself, Carlini et al. (2023) discussed memorization behavior in LLM regarding their factors of model size, continuation size, and context size, suggesting they are inter-correlated. Biderman et al. (2023a) also discussed memorization in LLM regarding their training phase and the overlap between different model sizes, showing the possibility of using smaller LLM to predict memorization of larger LLM.
Compared to the previous memorization studies, our research discussed memorized text and extended this scope to less and unmemorized text. Additionally, we also discussed both the encoding and decoding dynamics of generations. Finally, we analyze the prediction results of memorization.
@@ 3.1.1 Memorization Criteria
In our setting, we prompt the model with a specific context length and use greedy decoding to generate the continuation tokens based on that previous context. Then, we compare the predicted continuations with the actual continuations and calculate a memorization score using the following equation:
@@
In the above equation, $X=\{x_{0}\cdots x_{n}\}$ is the sequence of predicted continuation tokens, and $Y=\{y_{0}\cdots y_{n}\}$ is the sequence of true continuation tokens. $L(Y)$ means the length of the true continuation tokens. This score is the memorization score of the sequence $Y$ under context tokens $C=\{c_{0}\cdots c_{n}\}$.
From this equation, we can see that if $M(X,Y)==1$ means that the sequence $Y$ is completely memorized under the context sequence $C$, which we call this sentence K-extractable. Similarly, we can also say that the sequence $Y$ is completely not memorzied if $M(X,y)==0$ In addition to this token level-based accuracy, we collected all sentences and divided them into different ranges in our experiment.
@@ 3.1.2 Prediction Criteria
In this study, we also perform the prediction of memorization. We use two criteria to evaluate prediction performance, e.g., Token-Level Accuracy and Full Accuracy.
We define the prediction at a specific token at idx $i$ as $\hat{X}=\{\hat{x_{0}},\hat{x_{1}}\dots\hat{x_{n}}\}$ where each prediction $\hat{x_{n}}$ is a binary label indicating the token at this index is a memorized token or not. Similarly, the gold result is denoted as $\hat{Y}=\{\hat{y_{0}},\hat{y_{0}}\dots\hat{y_{n}}\}$ where is $\hat{y}$ is the golden binary label
@@ 3.2 Model Setting
We use the Pythia model Biderman et al. (2023b) to analyze the memorization behavior as Pythia provides LLMs trained across various sizes (70m to 12b) with the same training order using open-sourced Pile Gao et al. (2020) corpora. This provides the benefits of analyzing memorization behavior with a more stable model choice and experiment stability.
Specifically, we investigate the model size of [70m, 160m, 410m, 1b, 2.8b, 6.9b, 12b] where $m$ stands for million and $b$ stands for billion. We choose the model trained on the deduplicated pre-train data version to avoid the effect of duplicated sentences, as previous research reports that the chance to be memorized grows exponentially with the number of duplicates. . [^1] For more details, please refer to https://github.com/EleutherAI/pythia
@@ 4.1 Memorization Factors
This section discusses how the number of memorized sentences grows with different factors. We collect the number of sentences memorized or unmemorized across different extend under different model sizes. The results are plotted in Figure 2.
@@ 4.1.1 The Factor of Model Size
In this experiment, we discuss how model size affects the number of memorized and unmemorized contexts. From the Figure 2 (a), we can see that
  • We can see that sentences with low memorization scores are significantly higher than sentences with high memorization scores. This shows that even though we can find memorization in the LLM, most of the contents they are trained in are not memorized.
  • Among sentences with different memorization scores, the memorized sentence increases faster than others. This indicates that LLM tends to memorize sentences more than to remember a piece of it fully. We can also see that the number of sentences with low memorization scores, like 0 and 0.1, decreases with the increase of model size, showing that the LLM tends to memorize the corpora with the increase of model.
  • Additionally, we can see that the sentences that are being memorized or memorized do not scale linearly with the increase in model size. For example, we can see that for fully memorized sentences, the increase in sentence numbers is obvious from 70m to 2.8b compared to the increase from 2.8b to 12b. Similarly, the same decreasing trend can be seen from the figure for the memorized sentence. This shows that even with enough model size, the LLM may still be unable to memorize the whole corpora.
@@ 4.1.2 The Factor of Context Size and Complement Size
  • The decrease in memorized sentences with the complement size is not linear. We can see a noticeable decrease when changing the complement size from 32 to 48, where the number of memorized sentences is almost halved.
  • Additionally, from this figure, we can see that some sentences are firmly memorized, and even when we increase the complement size from 64 to 96, the number of memorized sentences does not change too much.
  • We can also see that the drop in the number of memorized sentences is more significant in the larger model. This suggests that the increase in model size significantly helps to memorize short context compared to longer content.
  • The increase of memorized sentences with the context size is also not linear. The increase in context leads to an above-linear increase in the number of memorized sentences. This suggests that with more information on the context, the more LLM can generate related continuations.
@@ 4.2 Memorization Transition Across Model Size
This section discusses how different memorized sentences transit between different model sizes. For example, when trained with a larger model, we study how many unmemorized sentences in a small-size model will transition to memorized or low-memorized sentences. We plot the transition matrix in Figure 3. This figure shows the transition from the 410m size model to the 2.8b size and 2.8b size to the 12b size model. We assigned sentences with memorization scores in the range of 0.2 with labels of very low, low, medium, high, and very high. From the Figure 3, we can conclude the following:
  • Most sentences will stay at their previous state when trained in a larger model. We can tell this by looking at the diagonal of those two transition matrices. Most sentences will stay in their original states.
  • Additionally, highly memorized sentences are harder to transit compared to low memorized sentences. Especially for highly memorized sentences, more $94\%$ of sentences will remain being memorized compared to low memorized sentences or unmemorized sentences.
  • We can also see with the increase in the model size, the sentence is more inclined to stay in its original state. Figure b shows the transition from the 2.8b model to the 12b model; we can tell that the transition probability at the diagonal is higher than the 410b to 2.8b transition matrix counterpart. This suggests that the memorized or unmemorized sentences tend to be fixed with the increase in model size.
  • We can also see that even with a highly memorized sentence, there is still a little chance that the sentence will be forgotten or poorly memorized. The sentence in low memorized sentence tends to be highly memorized.
@@ 4.3 Input Dyanamics
4.3.1 What triggers the memorized or unmemorized content?
Firstly, we begin with the input level, which is the pre-train corpora. We conduct this study by analyzing the n-gram statistics. First, we show the dynamics of 1-gram statistics in the
@@ 4.3.1 What triggers the memorized or unmemorized content?
From the statistics, we can see:
  • We can observe a very obvious boundary effect in this figure. At the index around 32 which is the first generated token, the frequency drops and rises for the memorized sentence (positive boundary effect) while rises and drops for the unmemorized sentences (negative boundary effect) while the frequency drop is more significant than the increase in the frequency of the memorized sentence.
  • For the half-memorized sentence, we see the frequency increase and drop around the index of 39 which is half the length of generated tokens. This shows that for the half-memorized sentence, the previous half is mostly memorized and the later half is mostly unmemorized meaning that the memorized tokens are distributed in a near continuous way rather than scatter in the generated tokens. We also notice the repetitive pattern of half-memorized sentences and unmemorized sentences which is that before the generation of the unmemorized token, we can see an increase and drop in frequency.
  • The positive boundary effect of a memorized sentence suggests that the memorized sentence is caused by the first several tokens having higher frequency showing the frequency of the initial tokens is more important than the following tokens, e.g remember the first several tokens make the following sentence to be memorized more easily. The negative boundary effect of the unmemorized sentences indicates the same results that the low frequency of the first several tokens makes the following sentence also easy to forget.
4.3.2 Model Size and the Role of Boundary Effect
@@
  • In the context and continuation, we can see the boundary effect even when we scale the calculation to the whole sentences. That is for memorized sentences, the frequency is higher in lower in the context but higher in the continuation while the unmemorized sentence shows the opposite.
  • However, the half-memorized sentence shows different results in that the frequency of continuation tokens is higher than the average frequency of context which is caused by the increase of frequency before reaching the first generated unmemorized tokens indicated in Figure 4.
  • In the specific boundary difference, we can see a decreasing trend for the positive boundary effect of memorized sentences while an increasing trend for the negative boundary effect of unmemorized sentences. These results suggest sentences with significant frequency increase are easy to memorize or forget.
@@
160m .708 .713 .744 55 69 .739 .837 .628 54 56 .12 -.22 -.78 3.3 -6.6
1b .713 .711 .752 55 69 .736 .832 .631 55 56 .11 -.27 -.76 2.8 -6.9
6.8b .721 .710 .759 .736 .829 .638 .09 -.15 -.74
12b .721 .720 .760 .736 .846 .626 .04 -.24 -.72
Table 1: Gram Statistics. $B$ means billion. $M$ means a million. We ignore the unit cardinal digit for the one-gram statistics as the unit digit is the same, which is $1$ for Context and Continuation while 00 for Freq Diff.

### v2

@@ Abstract
Memorization behavior of Large Language Models (LLMs) means LLMs can generate the same content used to train them. Though previous research has discussed memorization, the memorization of LLMs still lacks explanation, especially the cause of memorization and the dynamics of generating them. In this study, we comprehensively discussed memorization from various perspectives and extended the discussion to less and unmemorized content. Through various studies, we found that: (1) Through experiments, we revealed the relation of memorization between model size, continuation size, and context size. Further, we showed the dynamics of how unmemorized sentences transition to memorized sentences. (2) Through analysis of input, we found the frequency of tokens related to memorization and discovered the boundary effect when generating memorized or unmemorized sequences. (3) Through an analysis of embedding and decoding, we showed the LLM can generate paraphrased memorized sentences and sentences of different memorization scores clustered in the embedding space. Additionally, the entropy also shows a boundary effect when generating memorized or unmemorized content. (4)We trained a Transformer model to predict the memorization of different models, showing the possibility of predicting memorization and paraphrase memorization. Further analysis shows that the prediction of unmemorized sequences is easier than the prediction of memorized content, which can be explained by the significance of the boundary effect.
@@ 1 Introduction
Large language models (LLMs), spanning from BERT Devlin et al. (2019) to the recent strongest GPT-4 OpenAI et al. (2024), have not just revolutionized the Natural Language Processing (NLP) but nearly the whole artificial intelligence research field. Though their performances on various tasks are surprising, we still know little about the underlying mechanism of Large Language Models whose structures are based on neural models known as black-box Alain and Bengio (2018). Due to the enormous model size and pre-train data size, unprecedented phenomena in previous machine learning are also identified in the LLM, while in this research, we discuss the memorization.
Memorization Hartmann et al. (2023) in the context of LLMs means the LLM sometimes will verbatim output the same sentence recorded in their pre-training corpus. Memorization is a coin with two sides. On the one hand, utilizing such memorization, we can use LLMs as a knowledge base Petroni et al. (2019) that can be served for multiple uses. On the other hand, personal information contained in the pre-train corpora may also be elicited maliciously Yao et al. (2024). Though such a phenomenon has been utilized or studied since the merge of LLMs, detailed analysis of memorization is still under-explored.
@@ 2.1 Scaling Laws of LLM
In this study, our experiments span across various model sizes, thus relating to the research of Scaling Laws Kaplan et al. (2020); Abnar et al. (2021); Villalobos (2023), which suggests the LLM’s performance scales with the size of corpora, the parameter size and the computation required.
Those researches inspired LLMs to continuously scale those factors to gain higher performance like T5 Raffel et al. (2023), GPT-3 Brown et al. (2020), PaLM 2 Anil et al. (2023), etc. On the other hand, researchers also analyzed how LLM Scaling affects particular tasks rather than general performance. For example, Sun and Miceli-Barone (2024) discussed how LLM scaling affects both translation tasks and prompt injection attacks. Within those discussions, the LLM’s emergent abilities Wei et al. (2022) are also discovered, which means the LLM suddenly reaches high performance on previous low-performance tasks when the LLM reaches a specific model size such as digital addition Nye et al. (2021), reasoning Lampinen et al. (2022). Recent studies have questioned whether emergent abilities are just mirages caused by misuse of metrics Schaeffer et al. (2023) or whether emergent abilities are just context-learning Lu et al. (2023).
@@ 2.2 Memorization
Before LLM, a phenomenon close to memorization in real-world practices is known as over-fitting Tetko et al. (1995), which means a near-zero loss in the train set while bad performance in the test set. Zhang et al. (2017) described that the neural model could reach near-zero loss on the training set, which suggests the neural model perfectly memorized the relation between input and data label. However, memorization is different from over-fitting as LLM has good performance in the test set or the real world. Feldman (2020) has conducted a theoretical analysis explaining why memorization happens and is actually needed. They showed in a long-tail distribution where many categories only have very few samples. As the neural model cannot extract general features from the limited number of samples in those categories, the best choice for the neural model is just to memorize them and compare them with data in the test set.
In the era of LLM, different from classification models, the LLM can directly output their pre-training content that can be used from a knowledge graph Petroni et al. (2019); AlKhamissi et al. (2022) while also leads to the concern of data contamination Sainz et al. (2023) and privacy risk Yao et al. (2024). Regarding specific research on memorization behavior, Tirumala et al. (2022) discussed training dynamics of memorization, showing large models tend to memorize training samples more easily. Carlini et al. (2023) discussed memorization behavior in LLM regarding their factors of model size, continuation size, and context size, suggesting they are inter-correlated. Biderman et al. (2023a) also discussed memorization in LLM regarding their training phase and the overlap between different model sizes, showing the possibility of using smaller LLM to predict memorization of larger LLM.
Compared to the previous memorization studies, our research discussed memorized text and extended this scope to less and unmemorized text. Additionally, we also discussed both the encoding and decoding dynamics of generations. Finally, we trained a Transformer model to predict memorization and analyze the results.
@@ 3.1.1 Memorization Criteria
We prompt the LLM with a sequence of context tokens $C=\{c_{0}\cdots c_{n}\}$ whose length is defined as context length and use greedy decoding to generate the continuation tokens based on the prompted context. Then, we compare the predicted continuations with the actual continuations and calculate a memorization score using the following equation:
@@
In the above equation, $X=\{x_{0}\cdots x_{n}\}$ is the sequence of predicted continuation tokens, and $Y=\{y_{0}\cdots y_{n}\}$ is the true continuation tokens. $L(Y)$ means the length of the true continuation tokens. If $M(X,Y)==1$, this sequence is completely memorized under the context sequence, which called as K-extractable. Similarly, we also define the sequence $Y$ as completely not memorized if $M(X,y)==0$. Since the memorization score has high granularity, in some experiments, we also classify sentences into several memorization levels based on their memorization scores.
@@ 3.1.2 Prediction Criteria
In this study, we use two criteria to evaluate model performance in predicting memorization, e.g., Token-Level Accuracy and Full Accuracy.
We define the prediction for a sequence of generated tokens as $\hat{X}=\{\hat{x_{0}},\hat{x_{1}}\dots\hat{x_{n}}\}$ where each prediction $\hat{x_{n}}$ is a binary label indicating the token at this index is a memorized token or not. Similarly, the gold result is denoted as $\hat{Y}=\{\hat{y_{0}},\hat{y_{0}}\dots\hat{y_{n}}\}$ where is $\hat{y}$ is the golden binary label.
@@ 3.2 Model Setting
We use the Pythia model Biderman et al. (2023b) to analyze the memorization behavior as Pythia provides LLMs trained across various sizes with the same training order using open-sourced Pile Gao et al. (2020) corpora, which benefits the experiment and provides stability.
Specifically, we investigate the model size of [70m, 160m, 410m, 1b, 2.8b, 6.9b, 12b] where $m$ stands for million and $b$ stands for billion. We choose the model trained on the deduplicated pre-train data version to avoid the effect of duplicated sentences, as previous research reports that the chance to be memorized grows exponentially with the number of duplicates. [^1] For more details, please refer to https://github.com/EleutherAI/pythia
@@ 4.1 Memorization Factors
This section discusses how memorization is affected by different factors. We collect the number of sentences memorized across different extend under different model sizes. Then, we divide them into 10 sets with a memorization difference of 0.1 The results are plotted in Figure 2.
@@ 4.1.1 The Factor of Model Size
In this experiment, we discuss how model size affects the number of memorized and unmemorized contexts. From the Figure 2 (a), we can obtain that:
  • We can see that sentences with low memorization scores (0-0.3) are significantly higher than sentences with high memorization scores. This shows that even though we can find memorization in the LLM, most of the contents are still not memorized.
  • Among sentences with high memorization scores, the memorized sentence increases faster than others. This indicates that LLM tends to memorize sentences fully rather than memorizing a piece of it. We can also see that the number of sentences with low memorization scores, like 0 and 0.1, decreases with the increase of model size, showing that the LLM tends to memorize the corpora with the increase of model.
  • Additionally, we can see that the sentences that are being memorized or memorized do not increase or decrease linearly with the expansion in model size. For example, we can see that for fully memorized sentences, the increase in sentence numbers is obvious from 70m to 2.8b compared to the increase from 2.8b to 12b. Similarly, the same decreasing trend can be seen from this figure for the unmemorized sentence. This shows that even with enough model size, the LLM may still be unable to memorize the whole corpora.
@@ 4.1.2 The Factor of Context Size and Complement Size
  • The decrease in memorized sentences with the complement size is not linear. Especially we can see that some sentences are firmly memorized, and even when we increase the complement size from 64 to 96, the number of memorized sentences does not change too much compared to the change when the complement size changes from 32 to 48.
  • We can also see that the drop in the number of memorized sentences with the complement size is more significant in the larger model. This suggests that the increase in model size significantly helps to memorize short sequences rather than long sequences.
  • The increase of memorized sentences with the context size is also not linear. The increase in context leads to a near-exponential increase in the number of memorized sentences. This suggests that with more information in the context, the more LLM can generate corresponding continuations.
@@ 4.2 Memorization Transition Across Model Size
This section discusses how sentences with different memorization scores transit across model sizes. Specifically, when trained with a larger model, we study the mutual transition between sentences with different memorization scores. We assigned sentences with memorization scores in the range of 0.2 with labels of very low, low, medium, high, and very high. We plot the transition matrix in Figure 3, which shows the transition from the 410m size model to the 2.8b size and 2.8b size to the 12b size model.
From the Figure 3, we can conclude the following:
  • Most sentences will stay at their previous state even when trained in a larger model. We can tell this by looking at the diagonal of those two transition matrices. Additionally, the higher the memorization score, the harder the sentence will transition to another state. For highly memorized sentences, more than $90\%$ of sentences will remain being memorized compared to low memorized sentences.
  • We can also see with the increase in the model size, the sentence is more inclined to stay in its original state. In the transition from the 2.8b model to the 12b model, we can tell that the transition probability at the diagonal is higher than the 410m to 2.8b transition matrix counterpart. This suggests that the memorized or unmemorized sentences tend to be fixed with the increase in model size.
  • We also see that even for a highly memorized sentence, there is still a little chance that the sentence will transit to a low memorized state, suggesting some sentences may be memorized randomly rather than following a specific feature. Such randomness makes it possible for them to transit to a low memorized state when trained in a large model.
@@ 4.3 Input Dyanamics
4.3.1 Token Level Frequency Analysis
@@
Firstly, we begin with the input level, which is the pre-train corpora. We conduct this study by analyzing the n-gram statistics. First, we show the 1-gram statistics across steps in Figure 4, and we can see:
  • We can observe a very obvious boundary effect. At the index around 32, which is the first generated token, the frequency drops and rises for the memorized sentence (positive boundary effect) while rises and drops for the unmemorized sentences (negative boundary effect). Additionally, the negative boundary effect is more obvious than the positive boundary effect.
  • For the half-memorized sentence, we see the frequency increase and drop around the index of 39, which is half the length of generated tokens. This shows that for the half-memorized sentence, the previous half is mostly memorized, and the later half is mostly unmemorized, meaning that the memorized tokens are distributed in a near continuous way rather than scattered in the generated tokens. We also notice the repetitive pattern of half-memorized sentences and unmemorized sentences, which is the negative boundary effect around the index of 39.
  • The positive boundary effect of a memorized sentence suggests that the memorized sentence is caused by the first several tokens having a higher frequency, showing the frequency of the initial tokens is more important than the following tokens, e.g., remember the first several tokens make the following sentence to be memorized more easily. The negative boundary effect of the unmemorized sentences indicates the same results that the low frequency of the first several tokens makes the following sentence also easy to forget.
4.3.2 Sentence Level Frequency Analysis
@@
  • First, we can see that the frequency of one-gram is significantly higher than the frequency of two-gram. In the context and continuation, the average frequency is roughly 3.5 times higher than that of the two-gram frequency. Additionally, such boundary effect is consistent in the two-gram setting. However, the positive boundary effect is less obvious due to the decrease of frequency when computed by 2-gram. However, as the unit is billion, the actual frequency gap is still a million level.
  • In the context and continuation, in memorized sentences, the frequency is higher in lower in the context but higher in the continuation, while the unmemorized sentence shows the opposite. This suggests the boundary effect also exists in a broader sequence level, though less obvious.
  • However, the half-memorized sentence shows different results in that the frequency of continuation tokens is higher than the average frequency of context, which is caused by the increase of frequency before reaching the first generated unmemorized tokens indicated in Figure 4.
  • In the specific boundary frequency difference around the first generated token, we can see a decreasing trend for both positive and negative boundary effects. However, the decrease in the positive boundary effect makes it less significant, while the decrease in the negative boundary effect makes it more significant. These results suggest that the significance of the positive boundary corresponds to how easily the sentence can be memorized. In contrast, the significance of the negative boundary effect corresponds to how easily the sentence can not be memorized.
@@
160m 1.708 1.713 1.744 0.551 0.534 0.691 1.739 1.837 1.628 0.535 0.659 0.567 0.114 0.330 -0.939 0.033 0.101 -0.663
1b 1.713 1.711 1.752 0.558 0.552 0.697 1.736 1.832 1.631 0.509 0.682 0.564 0.103 0.270 -0.981 0.028 0.090 -0.696
6.8b 1.721 1.710 1.759 0.570 0.565 0.701 1.736 1.829 1.638 0.496 0.699 0.564 0.090 0.140 -0.963 0.027 0.085 -0.726
12b 1.721 1.720 1.760 0.572 0.569 0.702 1.736 1.846 1.626 0.493 0.704 0.563 0.039 0.237 -1.016 0.026 0.083 -0.732
Table 1: Gram Statistics. The frequency unit is billion. Boundary Freq Difference means we use the frequency of the generated token to subtract the frequency of the last token in the context.